Find the equation of the line that passes through the pair of points. Write your answer in slope-intercept form.

(−2, −4), (5, 3)

Answer by stanbon(75887)

Find the equation of the line that passes through the pair of points. Write your answer in slope-intercept form.

(−2, −4), (5, 3)
----
slope = (3--4)/(5--2) = 7/7 = 1
-----
Form:: y = mx + b
---
Solve for "b"::
3 = 1*5 + b
b = -2
-----
Equation:
y = x - 2
-------------
Cheers,
Stan H.
--------------